# RAGServe: Fast Quality-Aware RAG Systems with Configuration Adaptation

### v1

@@ Abstract
RAG (Retrieval Augmented Generation) allows LLMs (large language models) to generate better responses with external knowledge, but using more external knowledge often improves generation quality at the expense of response delay. Prior work either reduces the response delay (through better scheduling of RAG queries) or strives to maximize quality (which involves tuning the RAG workflow), but they fall short in optimizing the tradeoff between the delay and quality of RAG responses. This paper presents RAGServe, the first RAG system that jointly schedules queries and adapts the key RAG configurations of each query, such as the number of retrieved text chunks and synthesis methods, in order to balance quality optimization and response delay reduction. Using 4 popular RAG-QA datasets, we show that compared with the state-of-the-art RAG optimization schemes, RAGServe reduces the generation latency by 1.642.54×1.64-2.54$\times$1.64 - 2.54 × without sacrificing generation quality.
1 Introduction
@@
Retrieval-augmented generation (RAG) is a popular LLM inference technique that augments an LLM inference query with relevant text chunks, or “context”, retrieved from a large corpus. RAG systems, which include retrieval and LLM inference [^1] In the literature and industry lingo, RAG sometimes refers to the retrieval step, but in this work, RAG systems include both retrieval and LLM inference based on the retrieved texts, and we aim to optimize the whole pipeline., have found many use cases in QA tasks, personal assistants, chatbots, and LLM-powered search [55, 10]. While RAG can enhance the quality (accuracy and relevance) of LLM-generated responses [7, 47, 85, 80, 52], RAG queries are inherently slow as they need more compute and memory resources to process long input contexts to answer a query [6, 15, 39]. Thus, it is essential to balance high response quality and low response delays in RAG inference systems.
@@
Past research efforts have optimized RAG, regarding either response quality or response delay, but they fall short in optimizing the quality-delay tradeoffs of RAG. One line of prior work focuses on reducing response delay through better query scheduling (e.g., GPU allocation and inference batching) for RAG queries [40, 62, 66, 2], without changing the RAG queries themselves. An alternate line of work focuses on maximizing generation quality by tuning the configuration of RAG queries [73, 31, 67], but this is often done at the cost of longer response delay.
@@
The RAG configuration simultaneously affects generation quality and response delay. Unlike traditional data queries (e.g., SQL) which specify the inputs and operators, RAG queries are inherently under-specified as they contain only a short question written in natural language [51, 27, 57, 31]. The RAG system needs to first determine the RAG configuration, such as how many text chunks to retrieve (more in §2[69, 63, 73]. For instance, retrieving too many chunks for a simple RAG query may unnecessarily inflate delay without increasing quality.
@@
Moreover, multiple configuration knobs can be tuned to optimize delay-quality tradeoffs. For instance, besides how many chunks to retrieve, how to use them in the LLM’s input involves two design choices—should the chunks be processed by the LLM jointly, and should the text chunks be summarized first before being fed into the LLM together (and how long should a summary be). Some recent works also tune RAG configuration [31, 67], but they focus on tuning individual knobs and only maximize quality at the cost of higher delay.
@@
Finally, the impact of RAG configurations on quality-delay tradeoffs also varies significantly with queries. For example, to answer “In which country is the Kimbrough Memorial Stadium located?”, the RAG may retrieve and analyze one text chunk about the stadium. In contrast, to answer “Compare NVIDIA’s operating cost over the first three quarters of 2022 and identify the highest one”, the RAG may need multiple chunks, each containing the operating cost for a quarter, and process these chunks together, instead of reading them separately. Empirically, we show that picking RAG configuration per-query achieves $12-15\%$ higher quality and 2.53×2.5-3$\times$2.5 - 3 × lower delay than using any fixed configuration across all queries in a dataset (§5). Thus, RAG configurations should be adapted on a per-query basis. Yet, existing RAG systems, which hand-pick a static configuration offline based on a few example queries [36, 1, 21, 75], lose out on quality or response time.
@@
This paper presents RAGServe, the first RAG system that adapts multiple configuration knobs on a per-query basis and jointly makes configuration selections and scheduling decisions (i.e., which LLM inference in a batch) to optimize the delay-quality tradeoffs for RAG.
@@
At first, this would require solving a joint combinatorial problem for every query, which can be prohibitively expensive (§3). RAGServe tackles the challenge by a two-step approach.
@@
First, RAGServe prunes the massive configuration space for each received query to a smaller yet promising one that contains configurations that likely yield high-quality output for the query at hand. Specifically, RAGServe uses a separate LLM to estimate the query’s profile, including how many pieces of information is required to answer the query and whether joint reasoning is likely required across these pieces of information (more details in §4.1). The intuition of the query profiles is that they can effectively filter out undesirable RAG configurations. For the earlier query example “Compare NVIDIA’s operating cost over the first three quarters of 2022 and identify the highest one,” the estimated profile would suggest that it involves at least three separate pieces of information, so the number of chunks (one of the configuration knobs) should be at least three. It should be noted that the LLM-based profiler is an extra overhead in RAGServe, but fortunately, its input only contains the RAG query itself and the metadata of the RAG database, which are orders of magnitude shorter than the long contexts in RAG, so the estimation can be relatively fast, about 1/10 of the delay of the execution of the RAG query.
@@
Then, using the narrower configuration space, RAGServe reduces the RAG response delays by jointly deciding the per-query configuration and query scheduling (batching) based on available resources (§4.3). The insight is that with the pruned configuration space, the scheduler can pick the optimal configuration and batching decisions without exploring the original, large configuration space and the implication on quality.
@@
In short, RAGServe’s two-level design loosely decouples the problem into (1) pruning configuration space to a smaller yet promising range of configurations, which focuses solely on keeping the accuracy high, and (2) jointly optimizing configuration (within the narrowed range) and scheduling to optimize response delay by choosing configurations which best-fit into the GPU memory.
@@
Figure 1: Performance of RAGServe on the KG RAG FinSec [45] dataset compared to the baselines. Full results shown in §7.
@@
We evaluate RAGServe across four RAG datasets with diverse query profiles (e.g., reasoning vs. domain-specific QA). Figure 1 shows a preview of our results. Our key takeaways are as follows. When achieving the same or higher quality than the baselines, RAGServe reduces the response delay by 1.62.8×1.6-2.8$\times$1.6 - 2.8 × compared to the latest vLLM (a state-of-the-art serving engine), Parrot (the latest LLM query-scheduling method), as well as AdaptiveRAG (the latest RAG configuration-tuning method). RAGServe also achieves 1.84.5×1.8-4.5$\times$1.8 - 4.5 × higher throughput compared to these baselines when achieving the same response delay and same/higher quality.
2 RAG systems and configurations
@@
As an LLM often does not have domain-specific or up-to-date knowledge, LLM applications commonly employ RAG to supplement LLM inference with external knowledge to generate high-quality responses. Before processing queries, a RAG system organizes background documents by splitting them into chunks (each with a fixed number of tokens), embedding each chunk using models like Bert [12, 19], and storing the embeddings with the chunks in a vector database.
@@
    Retrieval: The RAG system retrieves one or more relevant context chunks from the database by comparing the query’s embedding, computed using the same embedding model during database indexing, with the stored embeddings.
@@
    Synthesis: After retrieving the relevant chunks, the RAG system combines these chunks and the RAG query to form an LLM call or multiple LLM calls to generate the answer for that query.
Retrieval is computationally lightweight and much faster than synthesis, so the response delay is typically dominated by the synthesis step [79].
@@
    How many chunks to retrieve (num_chunks): The number of context chunks directly affects the delay of the synthesis step, with more computation needed to process the longer sequences from using more chunks. In the meantime, retrieving too few chunks risks low response quality if the retrieved chunks do not contain enough useful information.
@@
    How to synthesize (synthesis_method): If more than one chunk is retrieved, then two underlying questions determine the choice of the synthesis method (illustrated in Figure 3). First, should the LLM read the chunks separately? If so, RAG uses the LLM to generate one answer for the query using each chunk separately and picks the output with the highest confidence, which is called map_rerank. This often incurs the least computation but can cause low quality if the useful information is scattered in different chunks, in which case the LLM should read the chunks jointly. The RAG system can choose to feed these chunks in the LLM input directly by concatenating them within a single prompt (referred to as stuff) or to create a shorter summary for each chunk first before feeding the summaries and the query into the LLM to generate the final response (referred to as map_reduce). stuff needs less computation than map_reduce, but risks degraded output quality for long inputs due to the lost-in-the-middle problem [42].
@@
Figure 2: The configuration knobs adapted by RAGServe are derived from key design choices of RAG systems.
While other tunable knobs exist (e.g., chunk size, embedding model, or LLM), tuning them requires costly reindexing of the data store or changes to the underlying embedding model or final LLM. In contrast, our configuration knobs only change the input of these two modules and thus require almost no code change to existing RAG systems.
@@
  • $\bullet$
    Response quality calculates the F1 score of the generated response against the ground truth. This metric is widely used in prior works [61, 63, 10].
  • $\bullet$
    Response delay measures the time elapsed from when the RAG system receives a RAG request to when it completes generating the response.
Figure 3: Illustration of different RAG synthesis methods, which have various LLM reasoning capabilities.
Next, we will show that these knobs need to be properly tuned on a per-query basis to achieve optimal tradeoff between quality and delay in §3.
3 Towards better quality-delay tradeoffs
Figure 4: Varying each RAG configuration knob leads to different quality-latency tradeoffs, and these tradeoffs differ across queries (Q1 in green, Q2 in blue, and Q3 in red).
Prior work on RAG either optimizes for lower delay or higher quality, but not the tradeoff between quality and delay. The first direction picks static configurations and focuses on reducing the delay by smart scheduling and resource allocation [40, 62, 66]. The second direction picks RAG configurations to maximize quality without regard to resource usage or delay [73, 31, 67]. For the first time, we explore the potential of optimizing the quality-delay tradeoffs for RAG.
@@ 3 Towards better quality-delay tradeoffs
To showcase this observation, we use three queries selected from Musique [68], a popular reasoning QA dataset (§7.1).
@@
We chose queries with different natural language complexity and reasoning, Q1 being relatively less complex than Q2 and Q3. Then, we adjust the value of each configuration knob in order to quantify each knob’s impact on the quality-delay tradeoffs in each of the three queries.
@@
For simple queries like Q1 (green), quality plateaus for more complex synthesis methods (stuff and map_reduce). Because it only needs a single piece of context, map_rerank that processes chunks in isolation suffices, whereas cross-chunk reasoning (stuff and map_reduce) adds undue delay (2×2$\times$2 ×) without improving quality.
In contrast, for queries such as Q2 (blue) that require cross-chunk reasoning, stuff and map_reduce provide significant quality improvements (35% increase) as they process all retrieved chunks jointly.
For even more complex queries, such as Q3 (red), which require even more reasoning and information (why Voyager 1 left has multiple reasons), methods like map_reduce improve quality (30% increase) by removing unnecessary text in the mapper phase, to help the LLM focus on the relevant content.
Figure 5: Per-query configuration can achieve significantly better quality-delay tradeoffs across queries compared to every fixed configuration choice.
Impact of the number of retrieved chunks:  Figure 4 (b) fixes the synthesis method (stuff) and shows the impact of the number of retrieved chunks (1 to 35) on quality and delay.
Simple queries, like Q1 (green), can often be answered using just one or two chunks (needs only birth county). For more complex queries, Q2 (blue) and Q3 (red), increasing the number of chunks from 1 to 15 improves the likelihood of retrieving all relevant context and improves quality.
Blindly retrieving more chunks than necessary risks diluting the relevance of actual important information, due to commonly known problems such as “lost-in-the-middle” [42]. In all three queries, retrieving more chunks beyond a point harms the quality (up to 20% drop) and unnecessarily inflates delay (up to 3×3$\times$3 ×). Hence we have a quality-delay tradeoff where increasing chunks up to a point helps quality but beyond that increases delay while degrading quality.
@@
Overall, we see that RAG queries naturally vary in complexity, requiring differing levels of reasoning between chunks and varying numbers of context chunks. More complex queries, which require more reasoning and context, benefit from increased LLM computation, which can come at the cost of increased delay. Adding more context chunks helps to a point beyond which it harms the output quality and delay.
Thus, adapting RAG configuration on a per-query basis is crucial.
Figure 5 uses queries from two datasets (Musique and QMSUM, see §7.1) and shows that picking the best configuration for each query (the best configuration is the one with the lowest delay that achieves less than 2% drop than the highest achievable quality) achieves superior quality-delay tradeoff than picking any static configuration for all queries. Specifically, choosing the configuration per-query allows up to 3$\times$ delay saving compared to static configurations which are the closest in quality. Every single static configuration choice that achieves comparable delay has at least a 10% quality drop.
In spite of the potential benefits, however, per-query configuration adaptation faces challenges that hinder their real-world adoption. Each RAG query comes in plain text with practically no RAG configurations associated with it. Moreover, the space of configurations grows exponentially with multiple knobs. For example, for a map_reduce configuration, having 30 values for num_chunks and 50 values for intermediate_length leads to 1500 configurations for a query. To enable per-query configuration selection, we cannot exhaustively profile all configurations for every query and choose the best.
Alternatively, if we profile periodically, we lose out on the potential configuration selection for each query, as variance in query profile leads to different quality-delay tradeoffs. Profiling cost is also prohibitively expensive as the LLM needs to be run with many synthesis methods, number of chunks etc., which require high GPU usage. Additionally, the delay of profiling can be $\sim$100$\times$ the inference delay due to multiple LLM calls during profiling. Online RAG queries have stringent requirements for GPU resource usage and end-to-end delay [62, 66]. This makes it hard to systematically decide what an optimal per-input configuration should be.
To truly achieve the benefit of per-query configuration adaptation, we need a smart system to drastically reduce the configuration space to useful ones, in a fast and cheap manner.
4 RAGServe: Enabling per-query configuration adaptation for RAG
We present RAGServe, a novel system for serving RAG queries with an emphasis on high generation quality and minimal delay. RAGServe is a RAG controller with two main components:

### v3

# METIS: Fast Quality-Aware RAG Systems with Configuration Adaptation

@@ Abstract.
RAG (Retrieval Augmented Generation) allows LLMs (large language models) to generate better responses with external knowledge, but using more external knowledge causes higher response delay. Prior work focuses either on reducing the response delay (e.g., better scheduling of RAG queries) or on maximizing quality (e.g., tuning the RAG workflow), but they fall short in systematically balancing the tradeoff between the delay and quality of RAG responses. To balance both quality and response delay, this paper presents METIS, the first RAG system that jointly schedules queries and adapts the key RAG configurations of each query, such as the number of retrieved text chunks and synthesis methods. Using four popular RAG-QA datasets, we show that compared to the state-of-the-art RAG optimization schemes, METIS reduces the generation latency by $1.64-2.54\times$ without sacrificing generation quality.
RAG systems, LLM inference, Scheduling
journalyear: 2025copyright: ccconference: ACM SIGOPS 31st Symposium on Operating Systems Principles; October 13–16, 2025; Seoul, Republic of Koreabooktitle: ACM SIGOPS 31st Symposium on Operating Systems Principles (SOSP ’25), October 13–16, 2025, Seoul, Republic of Koreadoi: 10.1145/3731569.3764855isbn: 979-8-4007-1870-0/2025/10ccs: Information systems Information systems applicationsccs: Computer systems organization Real-time system architecture
1. Introduction
@@
Retrieval-augmented generation (RAG) is a popular LLM inference technique that augments an LLM inference query with relevant text chunks, or “context”, retrieved from a large corpus. [^1] RAG vs. long-context models is an active field of research, with the industry widely deploying RAG for its task-focused model inference quality and better resource-sharing capabilities (rag_industry, ). RAG systems, which include retrieval and LLM inference, [^2] Though RAG sometimes refers to the retrieval step in this work, a RAG system includes both retrieval and LLM inference based on the retrieved texts, and we aim to optimize the whole pipeline., have found many use cases in QA tasks, personal assistants, chatbots, and LLM-powered search (ouyang2024contextaware, ; 10.1145/3626772.3657834, ). While RAG can enhance the quality (accuracy and relevance) of LLM-generated responses (balaguer2024ragvsfinetuningpipelines, ; mao-etal-2024-rag, ; zhu2024ragevalscenariospecificrag, ; zhao-etal-2024-optimizing, ; nguyen2024enhancingqadomain, ), RAG queries are inherently slow as they need more compute and memory resources to process the long input context to answer a query (acl_rag_tutorial, ; 10.1145/3637528.3671445, ; leng2024longcontextragperformance, ). Thus, it is essential to balance high response quality and low response delays in RAG inference systems.
@@
Past research efforts have optimized RAG, regarding either response quality or response delay, but they fall short in optimizing the quality-delay tradeoffs of RAG. RAG queries have an associated RAG configuration which describes how and how much data to input for the query (more in §2) (wang-etal-2024-searching, ; simon2024methodologyevaluatingragsystems, ; xie2024weknowragadaptiveapproachretrievalaugmented, ). One line of prior work focuses on reducing response delay through better query scheduling (e.g., GPU allocation and inference batching) for RAG queries (lin2024parrotefficientservingllmbased, ; shahout2024fastinferenceaugmentedlarge, ; tan2024teolaendtoendoptimizationllmbased, ; infercept, ; lin2025teleragefficientretrievalaugmentedgeneration, ), without adapting the RAG configuration themselves. An alternate line of work focuses on maximizing generation quality by tuning the configurations of RAG queries (xie2024weknowragadaptiveapproachretrievalaugmented, ; jeong-etal-2024-adaptive, ; tang2024mbaragbanditapproachadaptive, ), but this is often done at the cost of longer response delay.
@@
The RAG configuration simultaneously affects generation quality and response delay (e.g., retrieving too many chunks for a simple RAG query may unnecessarily inflate delay without increasing quality). Unlike traditional data queries (e.g., SQL) which specify the inputs and operators, RAG queries are inherently under-specified as they consist of a text query written in natural language (mombaerts2024metaknowledgeretrievalaugmented, ; guo2024lightragsimplefastretrievalaugmented, ; qian2024memoragmovingnextgenrag, ; jeong-etal-2024-adaptive, ) and do not directly specify the exact RAG configuration of its execution.
@@
Moreover, multiple configuration knobs can influence the delay-quality tradeoffs. For instance, besides how many chunks to retrieve, how to use them in the LLM’s input involves two design choices—should the chunks be processed by the LLM jointly, or should the chunks be summarized first before being fed into the LLM together (and how long should a summary be). Recent works also attempt to tune RAG configuration (jeong-etal-2024-adaptive, ; tang2024mbaragbanditapproachadaptive, ), but they focus on either tuning individual knobs or maximizing quality at the cost of higher delay. However, tuning configurations across multiple knobs quickly hits a prohibitive combinatorial space (more in §3) and requires optimizations to reduce the search cost.
@@
Finally, the impact of RAG configurations on quality-delay tradeoffs also varies significantly with queries. For example, to answer “In which country is the Kimbrough Memorial Stadium located?”, the RAG may retrieve and analyze one text chunk about the stadium. In contrast, to answer “Compare NVIDIA’s operating cost over the first three quarters of 2024 and identify the highest one”, the RAG may need multiple chunks, each containing the quarter’s operating cost, and process these chunks jointly, instead of reading them separately. The above examples illustrate queries differ in complexity (more in §4), leading to needing different configurations per-query for optimal quality-delay tradeoffs. Empirically, we show that picking RAG configuration per-query achieves $12-15\%$ higher quality and $2.5-3\times$ lower delay than using any fixed configuration across all queries in a dataset (§5). Thus, RAG configurations should be adapted on a per-query basis.
Yet, existing RAG systems, which hand-pick a static configuration offline based on a few example queries (kim2024autoragautomatedframeworkoptimization, ; rag_hyperparameter_tuning, ; autorag_hp, ; yao2024cacheblendfastlargelanguage, ), lose out on quality or response time.
@@
This paper presents METIS, the first RAG system that adapts multiple configuration knobs on a per-query basis and jointly makes configuration selections and scheduling decisions (i.e., which LLM inference in a batch) to optimize the delay-quality tradeoffs for RAG.
@@
As this would require solving a joint combinatorial problem for every query, which can be prohibitively expensive (§3), METIS tackles the challenge with a two-step approach.
@@
First, METIS prunes the massive configuration space for each received query to a smaller yet promising one that contains configurations that likely yield high-quality output for the given query. Specifically, METIS uses a separate LLM to estimate the query’s profile, including how many pieces of information are required to answer the query and whether joint reasoning is likely required across these pieces of information (more in §4.1). The intuition of the query profiles is that they can effectively filter out undesirable RAG configurations. For the earlier query example “Compare NVIDIA’s operating cost over the first three quarters of 2024 and identify the highest one,” the estimated profile would suggest that it involves at least three separate pieces of information, so the number of chunks (one of the configuration knobs) should be at least three. It should be noted that the LLM-based profiler is an extra overhead in METIS, but fortunately, its input only contains the RAG query itself and the metadata of the RAG database, which are orders of magnitude shorter than the long contexts in RAG, so the estimation can be relatively fast, about 1/10 of the delay of the execution of the RAG query.
@@
Using the narrowed configuration space, METIS reduces the RAG response delays by jointly deciding the per-query configuration and query scheduling based on available resources (§4.3). The insight is that within the pruned configuration space, the scheduler can make optimal configuration decisions without exploring the original, large configuration space and the implications on quality.
@@
In short, METIS’s two-level design loosely decouples the problem into (1) pruning configuration space to a smaller yet promising range of configurations, which focuses solely on keeping the accuracy high, and (2) jointly optimizing configuration (within the narrowed range) and scheduling to optimize response delay by choosing configurations which best-fit into the GPU memory.
@@
Figure 1. Performance of METIS on the KG RAG FinSec (kgrag, ) dataset compared to the baselines. Full results shown in §7.
@@
We evaluate METIS across four RAG datasets with diverse query profiles (e.g., reasoning vs. domain-specific QA). Figure 1 shows a preview of our results. Our key takeaways are as follows. When achieving the same or higher quality than the baselines, METIS reduces the response delay by $1.6-2.8\times$ compared to the latest vLLM (a state-of-the-art serving engine), Parrot (the latest LLM query-scheduling method), as well as AdaptiveRAG (the latest RAG configuration-tuning method). METIS also achieves $1.8-4.5\times$ higher throughput compared to these baselines when achieving the same response delay and same/higher quality.
The general concept of using LLMs to guide system tuning is not exactly new (298701, ; ong2025routellmlearningroutellms, ), but our key contribution lies in applying the concept to RAG systems, through joint scheduling with resource-aware configuration selection, leading to significantly better resource sharing (§4.2, §4.3). METIS is the first work which (a) shows the importance of tuning multiple RAG knobs; (b) profiles multiple knobs and adapts them simultaneously; and (c) is the first LLM system to introduce resource-quality tradeoff in its RAG decisions.
2. RAG systems and configurations
As an LLM often does not have domain-specific or up-to-date knowledge, LLM applications commonly employ RAG to supplement LLM inference with external knowledge to generate high-quality responses. Despite the growth of model context length, using RAG to pinpoint the relevant context is still significantly cheaper in terms of resource cost (GPU requirement), latency, and memory consumption (KV Cache size). For general-purpose QA pipelines, RAG is cost-efficient with retrieving targeted chunks based on semantic similarity to the query. Using LLMs with long-context documents in contrast has much higher GPU memory usage and delay (lin2025teleragefficientretrievalaugmentedgeneration, ; NEURIPS2024_a5d8aba2, ; li2024retrievalaugmentedgenerationlongcontext, ).
@@ 2. RAG systems and configurations
Before processing queries, a RAG system organizes background documents by splitting them into chunks (each with a fixed number of tokens), embedding each chunk using models like Bert (devlin2019bertpretrainingdeepbidirectional, ; mteb_benchmark, ), and storing the embeddings with the chunks in a vector database.
@@
    Retrieval: The RAG system retrieves one or more relevant context chunks from the database by comparing the query’s embedding, (using the same embedding model as for database indexing), with the stored embeddings.
@@
    Synthesis: After retrieving the relevant chunks, the RAG system combines these chunks and the RAG query to form a single/multiple LLM call(s) to generate the response.
Retrieval is computationally lightweight and much faster than synthesis (¿ $100\times)$, so the response delay is typically dominated by the synthesis step (ralmspec, ).
@@
    How many chunks to retrieve (num_chunks): The number of context chunks directly affects the delay of the synthesis step, with more computation needed to process the longer sequences with more chunks. In the meantime, retrieving too few chunks risks low response quality if the retrieved chunks do not contain enough useful information.
@@
    How to synthesize (synthesis_method): If the LLM should read the chunks separately, RAG uses the LLM to generate one answer for the query using each chunk separately and picks the output with the highest confidence, which is called map_rerank. This often incurs the least computation but can cause low quality if the useful information is scattered in different chunks, in which case the LLM should read the chunks jointly. The RAG system can feed these chunks in the LLM input directly by concatenating them within a single prompt (called stuff) or to create a shorter summary for each chunk first before feeding the summaries and the query into the LLM to generate the final response (called map_reduce). stuff needs less computation than map_reduce, but risks degraded output quality for long inputs due to the lost-in-the-middle problem (lost_in_the_middle, ).
@@
Figure 2. The configuration knobs adapted by METIS are derived from key design choices of RAG systems.
In this work, while we focus on universal RAG knobs which affect quality and delay common to all RAG systems, METIS can be extended to other tunable knobs (e.g., some RAG system may dynamically choose the embedding model, retrieval index or serving LLM). METIS’ design is extensible to any RAG configuration knob based on the query profile.
@@
Figure 3. Illustration of different RAG synthesis methods, which have various LLM reasoning capabilities.
Next, we show that these knobs need to be properly tuned on a per-query basis to achieve optimal tradeoff between quality and delay in §3.
3. Towards better quality-delay tradeoffs
Figure 4. Varying each RAG configuration knob leads to different quality-latency tradeoffs, and these tradeoffs differ across queries (Q1 in green, Q2 in blue, and Q3 in red).
Prior work on RAG either optimizes for lower delay or higher quality, i.e., the first picks static configurations and focuses on reducing the delay by smart scheduling and resource allocation (lin2024parrotefficientservingllmbased, ; shahout2024fastinferenceaugmentedlarge, ; tan2024teolaendtoendoptimizationllmbased, ) and the second picks RAG configurations to maximize quality without regard to resource usage or delay (xie2024weknowragadaptiveapproachretrievalaugmented, ; jeong-etal-2024-adaptive, ; tang2024mbaragbanditapproachadaptive, ). For the first time, we explore the potential of optimizing the quality-delay tradeoffs for RAG.
@@ 3. Towards better quality-delay tradeoffs
To showcase this observation, we use three queries from Musique (musique, ), a popular reasoning QA dataset (§7.1).
@@
We chose queries with different natural language complexity and reasoning, Q1 being relatively less complex than Q2 and Q3. Then, we adjust the value of each configuration knob in order to quantify each knob’s impact on the quality-delay tradeoffs in each of the queries.
@@
For simple queries like Q1 (green), quality plateaus for more complex synthesis methods (stuff and map_reduce). Because it only needs a single piece of context, map_rerank which processes chunks in isolation suffices, whereas cross-chunk reasoning (stuff and map_reduce) adds undue delay ($2\times$) without improving quality.
For queries such as Q2 (blue) that require cross-chunk reasoning, stuff and map_reduce provide significant quality improvements (35% increase) by processing chunks jointly.
For more complex queries, such as Q3 (red), which require even more reasoning and information (why Voyager 1 left has multiple reasons), methods like map_reduce improve quality (30% increase) by removing unnecessary text in the mapper phase, to help the LLM focus on the relevant content.
Impact of the number of retrieved chunks: Figure 4 (b) fixes the synthesis method (stuff) and shows the impact of the number of retrieved chunks (1-35) on quality and delay.
Simple queries, like Q1 (green), can often be answered using just one or two chunks (needs only birth county). For more complex queries, Q2 (blue) and Q3 (red), increasing the number of chunks (1-15) improves the likelihood of retrieving all relevant context and improves quality.
Blindly retrieving more chunks than necessary risks diluting the relevance of actual important information, due to commonly known problems such as “lost-in-the-middle” (lost_in_the_middle, ; hsieh2024rulerwhatsrealcontext, ). In all three queries, retrieving more chunks beyond a point harms the quality (up to 20% drop) and unnecessarily inflates delay (up to $3\times$). Hence we have a quality-delay tradeoff where increasing chunks up to a point helps quality but beyond that it increases delay while degrading quality.
@@
Overall, we see that RAG queries naturally vary in complexity, requiring differing levels of inter-chunk reasoning and varying numbers of context chunks. More complex queries, which require more reasoning and context, benefit from increased LLM computation, which can come at the cost of increased delay. Adding more context chunks helps to a point beyond which it harms the output quality and delay.
Thus, adapting RAG configuration on a per-query basis is crucial. Figures 2, 3, 4 illustrate tuning most popular RAG configuration knobs, however the tuning extends to more RAG configurations with richer tradeoff spaces (§4.2).
Figure 5. Per-query configuration can achieve significantly better quality-delay tradeoffs across queries compared to every fixed configuration choice.
[…]
In spite of the potential benefits, per-query configuration adaptation faces challenges that hinder their real-world adoption. Each RAG query comes in plain text with practically no associated RAG configurations. Moreover, the space of configurations grows exponentially with multiple knobs. For example, for a map_reduce configuration, with 30 values for num_chunks and 50 values for intermediate_length leads to 1500 configurations for a query. Exhaustively profiling all configurations per-query and choosing the best is infeasible.
[…]
To truly achieve the benefit of per-query configuration adaptation, we need a smart system to drastically reduce to a useful configuration space, in a fast and cheap manner.
Finally, in the emerging space of agentic and reasoning RAG, profiling and configuration adaptation remains an integral part due the latency-sensitive nature of these systems. Optimally choosing the configurations, the focus of METIS, remains crucial. We discuss this further in Section 9.
4. METIS: Enabling per-query configuration adaptation for RAG
We present METIS, a novel system for serving RAG queries focusing on high generation quality and minimal delay. METIS is a RAG controller (Figure 6) with two main components: